whats the perimeter of rectangle of 7inches and 3.2 inches?
Answer by jim_thompson5910(35256)

P = 2*(L+W)


P = 2*(7+3.2)


P = 2*(10.2)


P = 20.4


The perimeter is 20.4 inches